OpenStudy (anonymous):

Choose the alternative that best completes the statement or answers the question: cosx/tan^2x-cosx/sin^2x A) -cosx B) -sinx C) cosx D)sinx

OpenStudy (ybarrap):

$$ \cfrac{\cos x}{\tan^2x}-\cfrac{\cos x}{\sin^2x}\\ =\cfrac{\cos x\times\cos^2 x}{\sin^2x}-\cfrac{\cos x}{\sin^2x}\\ =\cfrac{\cos x\times\cos^2 x-\cos x}{\sin^2 x}\\ =\cfrac{\cos x(\cos^2x-1)}{\sin^2x}\\ =\cfrac{\cos x \sin^2x}{\sin^2x}\\ =\cos x $$ Make sense?

OpenStudy (anonymous):

isn't it a

OpenStudy (ybarrap):

Yes, you are right. Thanks for catching that! $$ \cfrac{\cos x}{\tan^2x}-\cfrac{\cos x}{\sin^2x}\\ =\cfrac{\cos x\times\cos^2 x}{\sin^2x}-\cfrac{\cos x}{\sin^2x}\\ =\cfrac{\cos x\times\cos^2 x-\cos x}{\sin^2 x}\\ =\cfrac{\cos x(\cos^2x-1)}{\sin^2x}\\ =-\cfrac{\cos x \sin^2x}{\sin^2x}\\ =-\cos x $$ Because $$ \sin^2x+cos^2x=1\\ \sin^2x=-cos^2x+1\\ \sin^2x=1-cos^2x\\ \implies \cos^2x-1=-\sin^2 x\\ $$